how many grams of Cu would have to react to produce 89.5 grams of Ag?


1
Expert's answer
2022-03-12T01:59:35-0500

It is unclear what reaction takes place, but it is reasonable to assume the general single displacement reaction:

Cu + 2Ag+ --> Cu2+ + 2Ag


"m(Cu)=89.5\ g(Ag)\times\frac{1\ mol(Ag)}{107.87\ g(Ag)}\times\frac{1\ mol(Cu)}{2\ mol(Ag)}\times\frac{63.55\ g(Cu)}{1\ mol(Cu)}=26.4\ g"


Answer: 26.4 g